the volume for a pyramid with a square base is V=1/3 wlh, where w is width, l is length and h is height. solve for w
thanks in advance

Answer by Adrish(44)

Solutions:
V=1/3wlh
Divide both sides by "lh"
V/lh=1/3 . wlh/lh
V/lh=w/3
Multiply both sides by 3
3V/lh=w